A curling rock is thrown and come to rest 40 m away 26 seconds later
a) what rate of acceleration is indicated?
b) what was the initial speed ?
c) sketch seprate graphs to show the position, speed and acceleration of the rock during those 26 seconds
1
Expert's answer
2019-10-24T09:23:59-0400

(a) Rate of acceleration is same for complete flight = "10\ m\ sec^{-2}"

(b)

"We\ know\ that ,\\R=\frac{u^2sin2\theta}{g}\\putting\ values,\\200=u^2sin\theta cos\theta\ \ \ .....(1)\\also,time\ of\ flight,T=\frac{2usin\theta}{g}\\or\\130=usin\theta\ .....(2)\\putting\ the \ (2)\ in\ (1),we\ get,\\ucos\theta=\frac{200}{130}\ \ ......(3)"

Dividing (2) from (3)

we get,

"tan\theta=84.5\\or\\\theta=89.322\degree\\Then,\\initial\ speed\ ,u=\frac{130}{sin89.322\degree}=130.01\ m\ sec^{-1}"